State clearly by what is meant by two events being statistically independent.
b. In 2015, a company estimated that 65% of its customers were female and, in that year, three customers
entered one of their shops. Assuming that these customers were not related or acquaintances, find the
probability that they were:
i. all female
ii. all male
iii. at least one male
c. In a certain factory which employs 10,000 men, 1% of all employees have a minor accident in a given
year. Of these, 40% had safety instructions whereas 90% of all employees had no instructions. What is the
probability of an employee being accident free:
i. Given that he had no safety instructions
ii. Given that he had safety instructions
1
Expert's answer
2020-03-24T06:35:10-0400

Linearly independent means one event has not any effect on the other event. In another way there is not any interception of the two event. In mathematically if A and B are linearly independent then,


"P(A \cap B)=P(A).P(B)\\"



b. Let event A=person is female & since entrance persons are not linearly dependent

i. "P(A)*P(A)*P(A)=0.65*0.65*0.65=0.2746"

ii "(1-P(A))*(1-P(A))*(1-P(A))=0.35*0.35*0.35=0.0429"

iii "P(at least one male)=1-P(all female)=1-0.2746=0.7254"


c.Let events,

A=has accident

B=has no accident

M=employees had instructions

N=employees hadn't instructions

"P(A)=0.01\\
P(B)=0.99\\
P(M|A)=0.4\\
P(N)=0.9\\
P(M)=0.1"


let ,

"P(N|A)=1-P(M|A)=0.6\\"

"P(N|B)=p;\\
P(N)=P(N|A).P(A)+P(N|B).P(B)\\
0.9 =0.6*0.01+p*0.99\\
p=0.903"

i. An employee being accident free given that he had no safety instructions "P(B|N),"

"P(B|N)=\frac{P(N|B).P(B)}{P(N)}\\
P(B|N)=\frac{0.903*0.99}{0.9}\\
P(B|N)=0.9933"


ii. An employee being accident free given that he had safety instructions P(B∣M),

"since, P(M|B)=1-P(N|B)\\
P(M|B)=1-0.903=0.097\\
P(B|M)=\frac{P(M|B).P(B)}{P(M)}\\
P(B|M)=\frac{0.097*0.99}{0.1}\\
P(B|M)=0.9603"